OpenStudy (anonymous):

Jack puts the letters of the word BUILDING on pieces of paper and puts the pieces in a bag. He draws two letters without looking.The first piece of paper is not replaced before the second piece is drawn. What is the probability that Jack draws an I twice. HELP!

OpenStudy (campbell_st):

what is the probability of drawing the 1st letter...?

OpenStudy (anonymous):

I'm not sure here are the answer choices A:1/28 B:1/32 C:3/28 D:11/28 i think its B

OpenStudy (campbell_st):

there are 2 I's and 8 letters so P(I) = 2/8 so if that I is not replaced what is the probability of drawing a 2nd I ?

OpenStudy (anonymous):

it would be 1/7 right

OpenStudy (campbell_st):

thats correct so P(I, I) = 2/8 x 1/7

OpenStudy (anonymous):

would that equal 22/56

OpenStudy (campbell_st):

nope \[\frac{2}{8} \times \frac{1}{7} = \frac{2 \times 1}{8 \times 7}\] and then simplify the fraction

OpenStudy (anonymous):

so it would be 1/28?

OpenStudy (campbell_st):

thats correct

OpenStudy (anonymous):

yesss, thank you also would the the answer to this question be 17.8, The probability that a dessert sold at a certain cafe contains chocolate is 74%. The probability that a dessert contains both chocolate and nuts is 24%. Find the probability that a randomly chosen chocolate dessert contains nuts. Round to the nearest tenth of a percent

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

yes... that makes sense to me....

OpenStudy (anonymous):

Okay thank you!